# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ Abstract
The interplay between short-range (SR) and long-range (LR) universal behaviors remains a fundamental topic in studying long-range interacting systems. The interaction between LR coupling and the Berezinskii-Kosterlitz-Thouless (BKT) mechanism introduces significant complexity, especially in the two-dimensional (2D) XY model, which is crucial for exploring low-dimensional phenomena and their implications for quantum computation. In the paper [arXiv:2404.08498], we investigated the 2D XY model with algebraically decaying interactions of the form $1/r^{2+\sigma}$. Our findings demonstrated continuous phase transitions into a ferromagnetic phase for $\sigma\leq 2$, characterized by the emergence of long-range order. In the ferromagnetic phase, for $\sigma<2$, we identified a power-law decaying correlation function attributed to the Goldstone mode, and for $\sigma=2$, logarithmic behaviors were observed. In this supplemental paper, we provide a comprehensive explanation of the detailed methodology, additional simulation results, and further insights into the above phenomena. It enhances our understanding of the crossover between SR and LR behaviors and has practical implications for experimental systems, such as Rydberg atom arrays.
@@ I Introduction
Long-range (LR) interactions are prevalent in various natural phenomena, such as gravitational interactions between celestial bodies and electromagnetic interactions between particles. A key characteristic of these interactions is that their strength decays with distance according to a power-law. Introducing LR interactions into complex many-body systems can significantly influence their phase transition properties, leading to a richer array of physical phenomena. The first study of LR interactions dates back to 1969, when Dyson investigated their effects in a one-dimensional (1D) Ising chain by introducing the interaction term $J_{ij}=|i-j|^{-(1+\sigma)}$ [1]. Dyson’s work revealed that, unlike the original Ising model, which exhibited no phase transitions, the LR interactions induce a second-order phase transition at finite temperatures for $10<\sigma<$ 1. This finding underscores the profound impact of LR interactions on phase transitions, demonstrating that they can lead to new and interesting physical behaviors in many-body systems.
@@
Upon this foundational work, the Ising model can be extended to the general $O(\mathcal{N})$ model with LR interactions, where the interaction decays as $r_{ij}^{-(d+\sigma)}$. In this context, $r_{ij}$ represents the distance between sites $i$ and $j$, and $d$ is the spatial dimension. This extension provides a broader framework for understanding the effects of LR interactions in various many-body systems.
@@
The long-range (LR) $O(\mathcal{N})$ model was first thoroughly studied by Fisher using a renormalization group (RG) approach [2]. He categorized the system’s behavior into three regimes based on the parameter $\sigma$: (1) $0<\sigma<\frac{d}{2}$ (classical regime): within this regime, the critical behavior of the system is governed by Gaussian fixed points; (2) $\frac{d}{2}<\sigma<2$ (non-classical regime): Gaussian fixed points cease to be stable, and the critical exponents vary with $\sigma$, with the critical exponent $\eta$ posited to precisely retain its value within the mean-field regime, namely $\eta=2-\sigma$; (3) $\sigma>2$ (short-range (SR) regime): in this interval, LR terms in the Hamiltonian become irrelevant, leading to the recovery of SR behavior. The regimes (1) and (3) have gained widespread acceptance, while regime (2) presents some issues. Firstly, the 1D LR Ising model undergoes a second-order phase transition for $10<\sigma<$ 1[1], a BKT transition at $\sigma=1$[3, 4], and exhibits no phase transition for $\sigma>1$, contradicting regime (2). Secondly, according to the regime (2), the exponent $\eta$ might exhibit a sudden jump from 0 to $\eta_{SR}$ at $\sigma=2$, where $\eta_{SR}$ denotes the $\eta$ value under short-range interactions, which is perplexing. Later, Sak proposed an alternative framework, known as the “Sak’s criterion”, where the boundary between the non-classical and short-range regimes shifts from 2 to $\sigma_{*}=2-\eta_{SR}$. In this revised framework, $\eta=2-\sigma$ is retained in the interval $1<\sigma<\sigma_{*}$, and $\eta=\eta_{SR}$ for $\sigma>\sigma_{*}$, thus simplifying $\eta$ to $\eta=\max(2-\sigma,\eta_{SR})$, eliminating the abrupt change in the exponent $\eta$. Sak’s criterion has been supported by a series of theoretical and numerical works [5, 6, 7, 8, 9, 10].
@@
However, later Monte Carlo (MC) results with higher precision and subsequent theoretical work suggest a different picture, where $\eta$ transitions smoothly from $2-\sigma$ near $\sigma=1$ to $\eta_{SR}$ at $\sigma=2$, and the threshold $\sigma_{*}$ returns to $2$ [11, 12]. Additionally, another MC study of a percolation model with LR probabilities seems to agree with this picture [13]. Later, another numerical study indicated the possible existence of double power in the correlation function near $\sigma=2-\eta_{SR}$, attributing this deviation to an underestimation of finite-size corrections [9].
@@
Overall, Sak’s criterion has achieved considerable recognition, but some controversies persist. Most numerical validations have been confined to the 2D LR Ising model, with limited exploration of LR $O(\mathcal{N})$ models for $\mathcal{N}\geqslant 2$. For instance, Ref.[14] studied the XY model on a 2D complex topology, which is expected to belong to the same universality class as the 2D LR XY model. Similarly, Ref.[15] investigated the 2D LR Heisenberg model. However, these studies did not focus on determining the threshold point $\sigma_{*}$, and their simulations were limited to relatively small system sizes, with a maximum length of only 256. Therefore, there is a pressing need for a more detailed numerical study of these models.
Additionally, the phase diagrams of the 2D LR XY model may exhibit unique characteristics worth exploring. The 2D XY model is known to undergo a BKT transition [3] with an exponent $\eta=0.25$, transitioning into a low-temperature phase with quasi-long-range order (QLRO). Previous studies have shown that the 2D LR XY model displays spontaneous magnetization at low temperatures for $\sigma<2$ [16]. This suggests that if Sak’s criterion holds for the 2D LR XY model, then within the interval $22-\eta_{SR}=1.75<\sigma<$ 2, the system should exhibit both a BKT transition and a low-temperature ferromagnetic phase. This scenario is particularly intriguing — recent work by Refs. [17, 18] has proposed a phase diagram where the LR XY system exhibits two phase transitions for $21.75<\sigma<$ 2: first a BKT transition into a QLRO phase, followed by a second-order transition into a ferromagnetic phase. Thus, a thorough and careful numerical investigation into whether the special phase diagram indeed exists is both necessary and meaningful.
@@
This paper is the accompanying paper of our previous work [19] to enhance further the understanding of the physics of the 2D LR XY model. In particular, we study the model via large-scale Monte Carlo (MC) simulations with the largest system size reaching $L=8192$. Instead of further refining the estimate of critical exponents in the non-classical regime, we focus on both the low-temperature and high-temperature properties of the model [19]. Our phase diagram for the LR XY model is shown in Figure 1. For $\sigma>2$, the system displays SR behavior: a BKT transition to QLRO for the XY model. For $\sigma<2$, the system exhibits a second-order phase transition, with critical exponents varying according to $\sigma$. In our research, we present a series of evidence that clearly indicates that the threshold between LR and SR universality is at $\sigma_{*}=2$ for the 2D LR XY model, which is inconsistent with Sak’s scenario and the proposed phase diagram in Ref. [17, 18].
@@
1. The second-moment correlation length ratio $\xi/L$ (will be defined in the later section) is a powerful tool for identifying the transition type and the critical point. Specifically, for a second-order transition, $\xi/L$ curves for different $L$s intersect at the critical point and diverge in the low-$T$ phases as $L$ increases. In contrast, for a BKT transition, $\xi/L$ curves for different $L$s converge to a smooth function in the QLRO phase. Our results show that for $\sigma\leq 2$, the scaling behavior of $\xi/L$ is consistent with the second-order transition, while for $\sigma>2$, typical BKT physics is observed.
@@
2. In the low-$T$ phases, we demonstrate that, for $\sigma\leq 2$, the system exhibits long-range ferromagnetic order in the thermodynamic limit, while for $\sigma>2$, the system enters a QLRO phase. In addition, the spontaneous symmetry breaking for $\sigma<2$ leads to Goldstone mode excitations, which give rise to a correlation function that decays algebraically to a constant, as $g(x)\sim x^{-\eta_{\ell}}+g_{0}$. Here, the anomalous dimension $\eta_{\ell}=2-\sigma$ is theoretically derived and confirmed numerically. Moreover, for the marginal case of $\sigma=2$, where theoretical derivation is absent, logarithmic behavior is observed in the low-$T$ phase.
@@
3. In the high-$T$ phases, as $T$ decreases to the critical point $T_{c}$, the growth behaviors of $\xi$ for $\sigma\leq 2$ and $\sigma>2$ are drastically different. For $\sigma>2$, typical BKT behavior is observed, $\xi\sim\exp(b/\sqrt{t})$, where $t=(T-T_{c})/T_{c}$ is the reduced temperature and $b$ is some non-universal constant. However, for $\sigma\leq 2$, the growth of $\xi$ increasingly deviates from the BKT behavior as $t\rightarrow 0$. It asymptotically follows the power-law divergences of a second-order transition as $\xi\sim t^{-\nu}$, where $\nu$ is the critical exponent of the correlation length.
4. Based on the previous knowledge of transition type, we carefully determine the critical temperature and critical exponents of the transitions for $\sigma\leq 2$. Our results do not support Sak’s scenario. Specifically, the anomalous dimension varies continuously in the regime $1<\sigma\leq 2$.
@@
In this auxiliary paper, we offer a more comprehensive explanation of our work, including the methodology, detailed analysis of the simulation results, and further insights to enhance our previous findings. In the following sections, we first describe the model and present the enhanced MC algorithm in Section II. The observables measured and their finite-size scaling analysis are also included. Sections III-VI then provide detailed results and analysis, including an overall explanation of the system at various $\sigma$, the low-temperature properties, high-temperature properties, and the critical point along with critical exponents. Finally, we summarize our findings in Section VII.
@@ II.1 Model
In this work, we investigate the LR XY model on a two-dimensional (2D) square lattice with periodic boundary conditions (PBC). The Hamiltonian of the model is given by,
@@
where $\bm{S}_{i}$ is a two-component unit vector corresponding to the XY spin at the $i$-th site, and $N=L\times L$ is the total number of spins with $L$ being the linear system size. The lattice spacing is set to unitary. The summation in Eq. (1) runs over all spin pairs, and the total number of interaction terms is $N_{b}=N(N-1)/2$. The spins interact with each other via an algebraically decaying ferromagnetic coupling,
$J_{ij}=\frac{J}{\left|\vec{r}_{i}-\vec{r}_{j}\right|^{2+\sigma}},$ (2)
where $J=1$ is the coupling constant, and $\vec{r}_{i}$ is the coordinate of the $i$-th site on the square lattice. In the $\sigma\rightarrow\infty$ limit, the model reduces to the nearest-neighbor (NN) XY model, which enters a quasi-long-range order phase at low-$T$ through the celebrated BKT transition [3]. In the $\sigma\rightarrow-2$ limit, this model becomes the XY model on a complete graph (CG), corresponding to the infinite-dimensional lattice with PBC in the thermodynamic limit [20, 21, 22]. In this work, we focus on the $\sigma>0$ regime, where the energy is strictly extensive, i.e., the energy density of the system remains finite in the thermodynamic limit.
To properly account for the long-range interactions across the periodic boundaries, the minimum-image convention is utilized [23, 24, 25]. The square lattice with PBC maps onto a torus, and spins interact via the shortest distance on the surface of the torus. Specifically, let $\vec{r}_{ij}$ being the displacement vector between sites $i$ and $j$, the $\alpha=x,y$ component of $\vec{r}_{ij}$ is calculated as,
$r_{ij}^{\alpha}=\begin{cases}r_{j}^{\alpha}-r_{i}^{\alpha}-L/2,&\text{if }r_{j}^{\alpha}-r_{i}^{\alpha}>+L/2\\ r_{j}^{\alpha}-r_{i}^{\alpha}+L/2,&\text{if }r_{j}^{\alpha}-r_{i}^{\alpha}<-L/2\\ r_{j}^{\alpha}-r_{i}^{\alpha},&\text{otherwise}\\ \end{cases}.$ (3)
In this model, the interaction range is not artificially truncated, and each spin interacts with all other $N-1$ spins.
The long-range interaction makes the average energy scale of the system depend on $\sigma$ and $L$, which can lead to a strong finite-size correction. To mitigate this correction and interpolate between the $\sigma\rightarrow\infty$ and $\sigma\rightarrow 0$ scenarios, we introduce a normalization constant $c(\sigma,N)$ to the Hamiltonian (1[26]. This normalization ensures that for all $\sigma>0$ and $L$, the average coupling strength for each spin equals that of the NN case,
$c(\sigma,N)^{-1}=\frac{1}{4J}\sum_{j=1}^{N-1}J_{0j}.$ (4)
Hence, the long-range interaction strength $J_{ij}$ becomes,
$J_{ij}=c(\sigma,N)\frac{J}{\left|\vec{r}_{i}-\vec{r}_{j}\right|^{2+\sigma}}.$ (5)
Although this normalization changes the averaged energy scale, it does not alter the algebraically decaying form of the interaction. Consequently, only the non-universal attributes of the phase transition, like the position of the critical points, are affected. The universal scaling properties at the critical point and the boundary between the LR and SR regimes remain unchanged. Hereafter, the interaction strength $J_{ij}$ refers to the normalized version defined in Eq. (5).
Another convention to mitigate the finite-size effects in long-range interacting systems is the Ewald summation technique, where the coupling constant $J_{ij}$ is modified to include the interactions of all periodic images of the system [8, 10]. This method is equivalent to our normalization treatment in the thermodynamic limit, only to differ in the finite-size correction terms to observables. For large-scale simulations, our method is more straightforward and asymptotically the same as the Ewald summation method in reducing the finite-size effect.

### v2

@@ Abstract
The two-dimensional (2D) XY model plays a crucial role in statistical and condensed matter physics. With the introduction of long-range interactions, the system exhibits a richer set of physical phenomena and a crossover between non-classical and short-range universality classes. In this work, we investigate the 2D XY model with algebraically decaying interactions $\sim 1/r^{2+\sigma}$, and provide a comprehensive numerical analysis of its thermodynamic properties. We demonstrate that for $\sigma\leq 2$, the system undergoes a second-order phase transition into a ferromagnetic phase characterized by the emergence of long-range order. In the low-temperature phase, due to the presence of the Goldstone mode, the correlation function saturates to a non-zero constant in the form of a power law for $\sigma<2$, with decaying exponent $2-\sigma$, and in the form of the inverse logarithm of distance for $\sigma=2$. Moreover, the critical points and exponents are also determined for various $\sigma$. We provide compelling evidence that the crossover between non-classical and short-range regimes occurs at $\sigma=2$. This work presents a detailed account of the simulation methodology, extensive numerical data, and new insights into the physics of long-range interacting systems.
@@ I Introduction
Long-range (LR) interactions are prevalent in various natural phenomena, such as gravitational interactions between celestial bodies and electromagnetic interactions between particles. A key characteristic of these interactions is that their strength decays with distance according to a power law. Introducing LR interactions into complex many-body systems can significantly influence their phase transition properties, leading to a richer array of physical phenomena. The first study of LR interactions dates back to 1969, when Dyson investigated their effects in a one-dimensional (1D) Ising chain by introducing the interaction term $J_{ij}=|i-j|^{-(1+\sigma)}$ [1]. Dyson’s work revealed that, unlike the original Ising model, which exhibited no phase transitions, the LR interactions induce a second-order phase transition at finite temperatures for $10<\sigma<$ 1. This finding highlights the profound impact of LR interactions on phase transitions, demonstrating that they can give rise to novel and intriguing physical behaviors in many-body systems.
@@
Building upon this foundational work, the Ising model can be extended to the general $O(n)$ spin model with LR interactions, where the interaction decays as $r_{ij}^{-(d+\sigma)}$. In this context, $r_{ij}$ represents the distance between site $i$ and $j$, and $d$ is the spatial dimension. This extension provides a broader framework for understanding the effects of LR interactions in various many-body systems.
@@
The LR $O(n)$ spin model was first thoroughly studied by Fisher using a renormalization group (RG) approach [2]. He categorized the system’s behavior into three regimes based on the parameter $\sigma$: (1) $0<\sigma<\frac{d}{2}$ (classical regime): within this regime, the critical behavior of the system is governed by the Gaussian fixed point; (2) $\frac{d}{2}<\sigma<2$ (nonclassical regime): Gaussian fixed points cease to be stable, and the critical exponents vary with $\sigma$, with the critical exponent $\eta$ posited to precisely retain its value within the mean-field regime, namely $\eta=2-\sigma$; (3) $\sigma>2$ (short-range (SR) regime): in this interval, LR terms in the Hamiltonian become irrelevant, leading to the recovery of SR behavior. Regimes (1) and (3) have gained widespread acceptance, whereas regime (2) presents some issues. According to the regime (2), the exponent $\eta$ might exhibit a sudden jump from 0 to $\eta_{\text{SR}}$ at $\sigma=2$, where $\eta_{\text{SR}}$ denotes the $\eta$ value under short-range interactions. Later, Sak proposed an alternative framework, known as Sak’s criterion, where the boundary between the nonclassical and short-range regimes shifts from 2 to $\sigma_{*}=2-\eta_{\text{SR}}$. In this revised framework, $\eta=2-\sigma$ is retained in the interval $1<\sigma<\sigma_{*}$, and $\eta=\eta_{\text{SR}}$ for $\sigma>\sigma_{*}$, thus simplifying $\eta$ to $\eta=\max(2-\sigma,\eta_{\text{SR}})$, eliminating the abrupt change in the exponent $\eta$. Sak’s criterion has been supported by a series of theoretical and numerical works [3, 4, 5, 6, 7, 8].
@@
However, later Monte Carlo (MC) results with higher precision and subsequent theoretical work suggest a different picture, where $\eta$ transitions smoothly from $2-\sigma$ near $\sigma=1$ to $\eta_{\text{SR}}$ at $\sigma=2$, and the threshold $\sigma_{*}$ returns to $2$ [9, 10]. Additionally, another MC study of a percolation model with LR probabilities appears to support this picture [11]. Later, another numerical study indicated the possible existence of a double power-law in the correlation function near $\sigma=2-\eta_{\text{SR}}$, attributing this deviation to an underestimation of finite-size corrections [7].
@@
Overall, Sak’s criterion has achieved considerable recognition, but controversies persist. Most numerical validations have been focused on the two-dimensional (2D) LR Ising model, with limited exploration of LR $O(n)$ spin models for $n\geqslant 2$. For instance, Ref. [12] studied the XY model on a 2D complex topology, which is expected to belong to the same universality class as the 2D LR XY model. Similarly, Ref. [13] investigated the 2D LR quantum Heisenberg model at finite temperatures. However, these studies did not focus on determining the threshold point $\sigma_{*}$, and their simulations were limited to relatively small system sizes, with a maximum length of only 256. Therefore, there is a pressing need for a more detailed numerical study of these models.
Apart from its implications for Sak’s criterion, the phase diagram of the 2D LR XY model exhibits novel characteristics. Although the Mermin–Wagner theorem forbids true long-range order (LRO) in a 2D XY model with finite-range coupling [14, 15], the system undergoes a BKT transition with an exponent $\eta_{\mathrm{SR}}=1/4$, entering a low-temperature (low-T) quasi-long-range ordered (QLRO) phase [16]. In contrast, in the mean-field regime $\sigma<\frac{d}{2}$, the XY model exhibits an ordinary second-order transition to a LRO phase. Thus, one would expect that in the nonclassical regime, there exists a point where the transition type changes drastically from BKT to second-order. The nature of this crossover regime remains a key open question. The Mermin–Wagner theorem applies under the condition that the second moment of the interaction $\sum_{\vec{r}}r^{2}|J(\vec{r})|$ is finite [14, 15]. For algebraically decaying interactions $J(r)\sim 1/r^{2+\sigma}$, this condition is violated when $\sigma\leq 2$ due to the divergence of the sum. This suggests that the theorem does not exclude the possibility of LRO for $\sigma\leq 2$. In 2001, Bruno provided a more stringent generalization of the Mermin-Wagner theorem for one- and two-dimensional LR XY and Heisenberg systems, stating that an LRO phase is allowed at sufficiently low temperatures for $\sigma<2$ but excluded for $\sigma\geq 2$ [17]. Moreover, a numerical study has shown that the 2D LR diluted XY model displays spontaneous magnetization at low temperatures for $\sigma<2$ [18]. This suggests that if Sak’s criterion holds for the 2D LR XY model, then within the interval $1.75<\sigma\leq 2$, the system should exhibit both a BKT transition and a low-T ferromagnetic phase. This scenario is particularly intriguing – recent works [19, 20] have proposed a phase diagram where the LR XY system exhibits two phase transitions for $21.75<\sigma<$ 2: first a BKT transition into a QLRO phase, followed by another transition into a ferromagnetic phase.
Figure 1: The phase diagram of the two-dimensional long-range XY model reveals distinct behaviors depending on the interaction range $\sigma$ and the inverse temperature $\beta=1/T$. In the short-range regime ($\sigma>2$), the system exhibits BKT transitions (brown line) into a QLRO phase. In the nonclassical regime ($1<\sigma\leq 2$), the system undergoes a continuous phase transition (red line) into a long-range ordered phase. Lastly, in the classical regime ($\sigma\leq 1$), the transition (purple lines) is governed by the Gaussian fix point. The symbols $\beta_{c}^{\text{CG}}$ and $\beta_{c}^{\text{NN}}$ denote the critical inverse temperatures for the complete-graph and NN interaction limits, respectively.
@@
This paper provides a comprehensive study of the 2D LR XY model, offering new insights into its critical behavior and phase diagram. We perform large-scale Monte Carlo (MC) simulations with the largest linear system size up to $L=8192$. Instead of further refining the estimate of critical exponents in the nonclassical regime, we focus on both the low- and high-temperature (high-T) properties of the model, and a brief summary of the main results is given in Ref. [21]. The phase diagram for the 2D LR XY model is shown in Figure 1. For $\sigma>2$, the system displays SR behavior: a BKT transition to QLRO for the XY model. For $\sigma\leq 2$, the system exhibits a second-order phase transition, with critical exponents varying according to $\sigma$. In this work, we present a series of evidence that strongly indicates that the threshold between LR and SR universality is at $\sigma_{*}=2$ for the 2D LR XY model, which is inconsistent with Sak’s scenario and the proposed phase diagram in Ref. [19, 20].
@@
1. The second-moment correlation length ratio $\xi/L$ (will be defined in the next section) is a powerful tool for identifying the transition type and the critical point. Specifically, for a second-order transition, $\xi/L$ curves for different $L$s intersect at the critical point and diverge in the low-T phases as $L$ increases. In contrast, for a BKT transition, $\xi/L$ curves for different $L$s converge to a universal function in the QLRO phase. Our results show that for $\sigma\leq 2$, the scaling behavior of $\xi/L$ is consistent with the second-order transition, while for $\sigma>2$, typical BKT physics is observed.
@@
2. In the whole low-T phases, we demonstrate that, for $\sigma\leq 2$, the system exhibits long-range ferromagnetic order in the thermodynamic limit, while for $\sigma>2$, the system enters a QLRO phase. In addition, the spontaneous symmetry breaking for $\sigma<2$ leads to Goldstone mode excitations, which give rise to a correlation function that decays algebraically to a constant, as $g(r)\sim r^{-\eta_{\ell}}+g_{0}$. Here, the anomalous dimension $\eta_{\ell}=2-\sigma$ is theoretically derived and confirmed numerically. Moreover, for the marginal case of $\sigma=2$, where theoretical derivation is absent, logarithmic behavior is observed in the low-T phase, i.e. $g(r)\sim 1/\ln r+g_{0}$.
@@
3. In the high-T phases, as $T$ decreases to the critical point $T_{c}$, the growth behavior of $\xi$ for $\sigma\leq 2$ and $\sigma>2$ are drastically different. For $\sigma>2$, typical BKT behavior is observed, i.e., $\xi\sim\exp(b/\sqrt{t})$, where $t=(T-T_{c})/T_{c}$ is the reduced temperature and $b$ is some non-universal constant. However, for $\sigma\leq 2$, the growth of $\xi$ increasingly deviates from the BKT behavior as $t\rightarrow 0$. It asymptotically follows the power-law divergence characteristic of a second-order transition as $\xi\sim t^{-\nu}$, where $\nu$ is the critical exponent of the correlation length.
4. We carefully determine the critical temperature and critical exponents of the transitions for $\sigma\leq 2$. The anomalous magnetic dimensions $\eta$ and the correlation-length exponent $\nu$ are continuous functions of $\sigma$ as long as $\sigma\leq 2$.
@@
The rest of this paper is organized as follows. We first describe the model and present the enhanced MC algorithm in Section II. The observables measured and their finite-size scaling analysis are also included. Sections III-VI then present detailed results and analysis, including an overall explanation of the system at various $\sigma$, the low-T and high-T behaviors, and the critical exponents along the line of phase transitions. Finally, we summarize our findings in Section VII.
@@ II.1 Model
We investigate the LR XY model on a two-dimensional square lattice with periodic boundary conditions (PBC), with the Hamiltonian as,
@@
where $\bm{S}_{i}$ is a two-component unit vector corresponding to the XY spin at the $i$th site, and $N=L\times L$ is the total number of spins. The summation runs over all spin pairs, and the total number of interaction terms is $N(N-1)/2$. The spins interact with each other via an algebraically decaying ferromagnetic coupling,
$J_{ij}=\frac{1}{\left|\vec{r}_{i}-\vec{r}_{j}\right|^{2+\sigma}},$ (2)
where $\vec{r}_{i}$ is the coordinate of the $i$th site. In the $\sigma\rightarrow\infty$ limit, the model reduces to the nearest-neighbor (NN) XY model, which enters a quasi-long-range order phase at low temperature through the celebrated BKT transition [16]. In the $\sigma\rightarrow-2$ limit, this model becomes the XY model on a complete graph (CG), which corresponds to the infinite-dimensional lattice with PBC in the thermodynamic limit [22, 23, 24]. In this work, we focus on the $\sigma>0$ regime, where the energy is strictly extensive, i.e., the energy density of the system remains finite in the thermodynamic limit.
To properly account for the long-range interactions across the periodic boundaries, the minimum-image convention is employed [25, 26, 27]. The square lattice with PBC maps onto a torus, and spins interact via the shortest distance on the surface of the torus. Specifically, the distance between any two lattice sites in the $x$ or $y$ direction does not exceed $L/2$.
In the simulation, we introduce a normalization constant $c(\sigma,N)$ which modifies Eq. (2) as
$J_{ij}=\frac{c(\sigma,N)}{\left|\vec{r}_{i}-\vec{r}_{j}\right|^{2+\sigma}},$ (3)
such that $\sum_{j}J_{ij}=4$. With such normalization, when $\sigma\to\infty$, the LR model reduces to the standard XY model with nearest neighboring interactions; when $\sigma\to-2$, $c(\sigma,N)=4/(N-1)$ also recovers the standard XY model on the finite complete graph. Another convention to deal with the long-range interactions in finite systems is the Ewald summation technique, where the coupling constant $J_{ij}$ is modified to include the interactions of all periodic images of the system [6, 8]. In the thermodynamic limit, this approach is equivalent to our normalization scheme, and the two only exhibit minor differences in finite systems, which quickly vanish as $L$ increases.